# QCRD: Quality-guided Contrastive Rationale Distillation for Large Language Models

### v1

@@ 4.1 Experimental setting
Datasets. To validate QCRD, we conducted extensive experiments on four widely-used benchmark datasets (see details in Appendix Table 1) across three different natural language processing tasks, including SVAMP Patel et al. (2021) for arithmetic word problem solving, CQA Talmor et al. (2018) for commonsense question answering, as well as e-SNLI Camburu et al. (2018) and ANLI Nie et al. (2019) for natural language inference. The rationales we used were generated by GPT-3.5-turbo https://platform.openai.com/docs/models and open-sourced by Hsieh et al. (2023).
Implementation details. Following the properties of CoT and the comparative experimental studies in Hsieh et al. (2023), our QCRD adopted T5-base (220M parameters) and T5-small (60M parameters) as the smaller (student) model, respectively. $\alpha_{1}$ and $\alpha_{2}$ were set to $0.5$, and $\alpha_{3}$ was initially set to $0.5$ and multiplied by $0.9$ in each iteration. We also set $\beta=0.2$ and $\delta=4$. We sampled the LLM’s output five times with the temperature parameter being $0.7$, and sampled a five-iteration-before model with the temperature parameter being $1.5$. We trained our models on one A100-80G GPU and the batchsize is 32.
Baselines. We followed the comparison methods in Hsieh et al. (2023). Specifically, three methods in learning task-specific models were compared, i.e., (1) Finetuning, which is the standard finetuning with the prevailing pretrain-then-finetune paradigm that finetunes a model with ground-truth labels via standard label supervision Howard and Ruder (2018); 2) Single-Task, where the student model is distilled to predict labels with the teacher model’s predicted labels; 3) DSS Hsieh et al. (2023), where the smaller model is distilled with both the predict labels and rationales of the LLM.
Figure 5: Comparisons with varying sizes of training datasets on T5-base model.
@@ 4.2 Experimental results
Experiments across four benchmarks. We conducted expriments across four benchmarks with two types of T5-model to evaluate the effectiveness of our proposed method. In Table 1, we summarized the experimental results of the T5-base model distilled by our method and the baselines individually on all the four datasets. Of note, in Single-Task, the rationale and label were combined into a single sequence, which was then treated as the target during the training Hsieh et al. (2023). One can readily see from Table 1 that our method achieved superior performance over the baselines on all the four datasets.
@@
Ours 69.0 63.64 54.00 90.15
In like manner, we performed our method and the baselines individually on the T5-small model, and their performance on all the four datasets was presented in Table 2. Our method consistently surpassed the baselines on all the four datasets as expected.
@@
Distillation with LLM labels. To evaluate the impact of label qualities on CoT distillation, without loss of generality, we conducted additional experiments on the three datasets (namely, CQA, ANLI1, and ESNLI) using the T5-base model distilled by our method and DSS. Instead of using ground truth labels, we employed the labels generated by GPT-3.5-Turbo to distill the student models. The results were presented in Table 3. On one hand, from the top of Table 3, it demonstrates the effectiveness of temperature sampling and self-consistency, which help denoise rationales and their corresponding labels. On the other hand, the results at the bottom of Table 3 indicate that our method outperforms DSS on CQA and ANLI1, even when utilizing the labels generated by the noise-affected LLM. It suggests that our method is more robust and can yield improved performance.
Table 3: CoT distillation results on the T5-base model.
ModelP CQA ANLI1 ESNLI
@@
Distillation with smaller datasets. In addition, to demonstrate the superiority of our method on smaller datasets, we compared the performance of DSS and our method using T5-base models on varying sizes of each of the four datasets. Fig. 5 illustrated that our method achieved better performance in most cases. It indicate the robustness and generality of our proposed QCRD.
Ablation study on our QCRD. Compared to previous related methods, the contrastive distillation in our QCRD introduces several key enhancements as follow. (1) The extension and denoising for positive knowledge (ED): we sample the output of the LLM and leverage the self-consistency to denoise rationales. (2) The distillation for negative knowledge (NK): we incorporate a self-play mechanism to generate low-quality rationales as negative samples. (3) The guidance of the Quality Judge (QJ): the use of a quality judgement discriminator helps assess rationales and optimize the training process. Additional experiments were so conducted on SVAMP to evaluate the effectiveness of each module, with the results being summarized in Table 4. The findings demonstrate that integrating more high-quality rationales significantly improves performance, while the inclusion of negative rationales proves effective. The discriminator mechanism plays a positive role by considering the quality of each rationale, and we further find that the results when using the Quality Judge are more stable.
Table 4: Ablation study on T5-base model.
@@ 5 Discussion
5.1 Different schemes for contrastive distillation
In sec. 3.3.2, we define the many-to-one distillation by taking the min loss for the positive rationales and max loss (after margin operation) for the negative rationales, which imposes a relatively weak constraint on rationale alignment. We here further discuss different schemes for the many-to-one distillation. (1) MaxMin: we compute the max loss for the positive rationales and min loss for the negative rationales. This scheme enforces the smaller model to learn hard rationale examples. (2) Sampling: we randomly choose a positive rationale and a negative rationale for each anchor. (3) Mean: we average the loss for all rationales. (4) Weighted mean (W-mean): we weight the loss with quality scores and then average the loss. The results of the T5-base model distilled by our method on SVAMP were presented in Table 5 with respect to the above different schemes. One can clearly see that MinMax achieved the best performance. Besides, the scheme of Mean had a negative impact on the results. The reason may be that the positive and negative rationales generated by temperature sampling may be of great difference, and especially for the positive rationales, enforcing the small model align with multi-target outputs is not suitable.
@@ 5.2 Influence of the sampling count
In the above experiments, we sampled the output of the LLM five times and the output of iteration-before model once. We further explore the influence of the sampling count. When fixing the sampling counts for the iteration-before model, the results of setting different sampling counts for the LLM on SVAMP were displayed in the top of Table 6. Moreover, when fixing the sampling counts for the LLM, the results on SVAMP were displayed in the bottom of Table 6 in terms of different numbers of generated negative samples. We found that sampling many negative rationales had an adverse impact on the performance, and the best performance was achieved when $k$ was 5. Note that when $k=1$, the performance of our method was still better than that of other related methods, again indicating the effectiveness of the use of negative rationales.
Table 6: Results of our method on SVAMP with different sampling counts.
@@
5.3 Rational analysis of negative rationales by self-play
Temperature sampling is a commonly used decoding strategy for LLMs’ generation process. In our method, we generated low-quality rationales by sampling the iteration-before smaller models. As displayed in Appendix Figure 2, we provided a case of the output rationale with different temperature settings for better illustration. The golden rationale can be obtained by greedy search; however, when the temperature was set to a higher value, the output rationales are likely various but inaccurate.
The results of the T5-base model distilled by our method on SVAMP were displayed in Table 7 in terms of different negative sampling temperature settings. It was observed that when no sampling was performed (i.e., $\tau=0$) or a lower temperature value was used (i.e., $\tau=0.7$), the smaller model exhibited relatively poorer performance and showed larger fluctuations in accuracy. The best results were achieved when the temperature $\tau$ was set to $1.5$. The reason for this can be attributed to the fact that when the model approaches convergence, the output rationales with lower temperature values tend to be similar to the golden ones. Considering these similar outputs as negative samples can lead to detrimental effects. All in all, we sampled the output rationales of iteration-before models several times and utilized these generated low-quality rationales as negative samples to enforce the small models away from these distributions and boost the performance.
Table 7: Results of our method on SVAMP with different negative sampling temperature settings.
$\tau$ 0 0.7 1 1.5 2
T5-base 65.0 64.5 67.5 69.0 67.5
5.4 The quality assessment for the output rationales
We assessed the quality of CoT examples using GPT-3.5-turbo. Inspired by the ranking model, we prompted GPT-3.5-turbo to rank the rationales generated by both DSS and our QCRD, rather than providing scores based on the quality of the rationales. This is easier for the LLM and allows for a more straightforward comparison. The prompt provided to GPT-3.5-turbo was presented in Table 8. To evaluate the rationales, we randomly selected $50$ examples from each of the four datasets and asked GPT-3.5-turbo to determine rationales of which our method was better. We then aggregated the counts of "DSS is better," "Both are good," and "QCRD is better," as shown in Table 9. From the results, we observed that on SVAMP, CQA, and ESNLI, the model trained using our method generated better rationales than using DSS. However, on ANLI1, the model trained using DSS exhibited slightly better performance.
Table 8: The prompt for GPT-3.5-turbo to judge rationales.
The prompt for GPT-3.5-turbo
Table 9: The quality assessment results on T5-base model for different sampling temperature settings, where three numbers represent the counts of "DSS is better", "Both are good", and "QCRD is better", respectively.
@@ 6 Conclusion
The knowledge distillation of CoT rationales from LLMs into smaller reasoners using a multi-task framework has been empirically shown to enhance performance. Building upon this framework, this paper introduces general CoT distillation method, incorporating a contrastive learning perspective that considers both positive and negative knowledge. To generate positive and negative rationales, we propose an innovative approach that combines temperature sampling, the self-consistency of LLMs, and the self-play of the small reasoners themselves. Additionally, we develop a many-to-one contrastive distillation loss for better knowledge learning, where an online-update discriminator is used to judge the qualities of rationales and assign weights for optimizing the training process across the whole datasets. Extensive experiments conducted on multiple reasoning tasks demonstrate the superiority of our method over the previous ones.
@@ Appendix A Appendix
A.1 Iteration-before-models for negative rationale generators
In this paper, we dynamically generated negative rationales using iteration-before-models through online temperature sampling. As depicted in Fig. 1, to select a j-iteration-before-model for the negative rationale generator, we need to save a minimum of $j$ checkpoints for the model. This allows us to train the model end-to-end instead of using a multi-turn approach.
Figure 1: A case of the j-iteration-before-model for negative rationale generator.
A.2 Details about the Quality Judge
We incorporated a discriminator into our training process to assess the quality of rationales and assign corresponding weights to the losses. To construct the discriminator, we leveraged the encoder of the T5-base model along with two additional linear layers to compute the quality score. Prior to training, the discriminator needs to be pretrained using the output rationales generated by LLMs with applying data augmentations to the negative rationales. Specifically, we employed word replacement with the assistance of StanfordNLP Zeman et al. (2018). Additionally, throughout the training process, we dynamically updated the discriminator using a decayed learning rate for better learning the negative cases from iteration-before-models.
Figure 2: A case study of the output rationale on SVAMP with different temperature sets.

### v2

@@ 4.1 Experimental setting
Datasets. We conducted extensive experiments on four widely-used benchmark datasets (see details in Appendix Table 1) across three different natural language processing tasks, including SVAMP Patel et al. (2021) for arithmetic word problem solving, CQA Talmor et al. (2018) for commonsense question answering, as well as e-SNLI Camburu et al. (2018) and ANLI Nie et al. (2019) for natural language inference. The rationales we used were generated by GPT-3.5-turbo [^1] https://platform.openai.com/docs/models and an opened code source by Hsieh et al. (2023) was referred.
Implementation details. Following the properties of CoT and comparative experimental studies in Hsieh et al. (2023), our QCRD adopted T5-base (220M parameters) and T5-small (60M parameters) as the student model, respectively. $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ were set to $0.5$ empirically. $\alpha_{3}$ was multiplied by $0.9$ per iteration. We set $\beta=0.2$ and $\delta=3$. We sampled the LLM’s output $5$ times with the temperature being $0.7$, and sampled $5$-iteration-before models with the temperature being $1.5$. The batchsize was 8 and learning rate was 1e-5. We trained our models with 10000 max steps on one A100-80G about 13 hours for T5-base and 8.5 hours for T5-small. The reported metric was accuracy.
Baselines. Four methods in learning task-specific models were compared, i.e., (1) Finetuning, which is the standard finetuning with the prevailing pretrain-then-finetune paradigm that finetunes a model with ground-truth labels via standard label supervision Howard and Ruder (2018); (2) Single-Task, where student models are distilled to predict labels with the teacher model’s predicted labels; (3) DSS Hsieh et al. (2023), where student models are distilled with both the predict labels and rationales of the LLM; (4) Mutual information (MI) Chen et al. (2024b), which is based on DSS and applies an additional task to maximizing the mutual information between prediction labels and rationales.
Figure 4: Comparisons with varying sizes of training datasets on the T5-base model for four benchmarks.
@@ 4.2 Experimental results
Experiments across four benchmarks. We conducted experiments across four benchmarks with two types of T5-model to evaluate the effectiveness of our proposed method. In the top of Table 1, we summarized the experimental results of the T5-base model distilled by our method and the baselines individually on all the four datasets. Of note, in Single-Task, the rationale and label were combined into a single sequence, which was then treated as the target during the training Hsieh et al. (2023). It is clear that our method outperformed the baselines on most datasets, particularly when compared to the baseline DSS.
@@
MI 67.5 63.50 54.20 90.15
Ours 69.0 63.64 54.00 90.26
In like manner, we performed our method and the baselines individually on the T5-small model, and their performance on all the four datasets was presented in Table 2. Our method consistently surpassed the baselines on all the four datasets.
@@
Distillation with LLM labels. To evaluate the impact of label qualities on CoT distillation, without loss of generality, we conducted additional experiments on the three datasets (namely, CQA, ANLI1, and ESNLI) using the T5-base model distilled by our method and DSS. Instead of using ground truth labels, we employed the labels generated by GPT-3.5-turbo to distill student models. The results were presented in Table 3. On one hand, from the top of Table 3, it demonstrates the effectiveness of temperature sampling and self-consistency (SC), which help denoise rationales and their corresponding labels. On the other hand, the results at the bottom of Table 3 indicate that our method outperformed DSS on CQA and ANLI1, even when utilizing labels generated by the LLM.
Table 3: CoT distillation results on the T5-base model using predicted labels (noisy labels) from the LLM.
Model CQA ANLI1 ESNLI
@@
Distillation with smaller datasets. In addition, to demonstrate the superiority of our method on smaller datasets, we compared the performance of DSS and our method using T5-base models on varying sizes of each of the four datasets. Fig. 4 illustrates that our method achieved better performance in most cases. It indicates the robustness and generality of our proposed QCRD.
Ablation study on QCRD. Compared to previous related methods, the contrastive distillation in our QCRD introduces several key enhancements as follow. (1) The extension and denoising for positive knowledge (ED): we sample the outputs of the LLM and leverage the self-consistency to denoise rationales. (2) The distillation for negative knowledge (NK): we incorporate a self-supervised mechanism to generate low-quality rationales as negative rationales. (3) The guidance of the Quality Judge (QJ): the use of discriminator helps assess rationales and optimize the training process. Additional experiments were so conducted on SVAMP to evaluate the effectiveness of each module, with the results being summarized in Table 4. The findings demonstrated that integrating more high-quality rationales significantly improved performance, while the inclusion of negative rationales proved effective. The discriminator mechanism played a positive role by considering the quality of each rationale, and we further found that the results when using the Quality Judge were more stable.
Table 4: Ablation study on T5-base model, where ED denotes positive knowledge extension and denoising, NK denotes negative knowledge, and QJ denotes using of Quality Judge.
@@ 5 Discussion
5.1 Different contrastive distillation schemes
In sec. 3.3.2, We defined the many-to-one distillation by taking the min loss for positive rationales and max loss for negative rationales (i.e., MinMax), which imposes a relatively weak constraint on rationale alignment. We further discuss different schemes for the many-to-one distillation. (1) MaxMin: we compute the max loss for positive rationales and min loss for negative rationales. This scheme enforces the smaller model to learn hard rationale examples. (2) Sampling: we randomly choose a positive rationale and a negative rationale for each input. (3) Mean: we average the loss for all rationales. (4) Weighted mean (W-mean): we weight the loss with quality scores and then average the loss. The results of the T5-base model distilled by our method on SVAMP were presented in Table 5 with respect to the above different schemes. One can clearly see that the MinMax achieved the best performance. Besides, the Mean scheme had a negative impact on the results. The reason may be that enforcing small models align with multi-target rationales of differences is not suitable, especially for positive knowledge.
@@ 5.2 Influence of the sampling count
In the above experiments, we sampled the output of the LLM five times and the output of iteration-before model once. We further explore the influence of the sampling count. When fixing the sampling counts for iteration-before models, results of setting different sampling counts for the LLM on SVAMP were displayed in the top of Table 6. Moreover, when fixing the sampling counts for the LLM, results on SVAMP were displayed in the bottom of Table 6 in terms of different numbers of generated negative samples. We found that sampling many negative rationales had an adverse impact on the performance, and the best performance was achieved when $k$ was 1. Note that when $m=1$, the performance of our method was still better than that of other related methods, again indicating the effectiveness of negative rationales.
Table 6: Results of our method on SVAMP with different sampling counts, i.e., the sampling count $m$ for positive rationales and $k$ for negative rationales.
@@
5.3 Rationality for negative knowledge
Temperature sampling is a commonly used decoding strategy for LLMs’ generation process. By adjusting the temperature $\tau$, we can modify the probability distribution of each word before sampling. The higher the temperature is, the smaller the difference in the probability distribution of LLM’s outputs becomes, increasing the chance of sampling words with lower probabilities. In Fig 5, we provided a case of output rationales from the trained T5-base model with different temperature settings for illustrative visualization. It validated the rationality that we generated negative rationales by sampling the iteration-before smaller models with a high temperature value. We further explored the influence of the negative sampling temperature on model performance in Appendix A.4.
Figure 5: A case study of output rationales from the T5-base model on SVAMP with different temperatures.
5.4 Assessment for generated rationales
We assessed the qualities of CoT examples using GPT-3.5-turbo. Inspired by the ranking model, we prompted GPT-3.5-turbo to rank the rationales generated by both DSS and our QCRD, rather than providing scores based on the qualities of the rationales. This is easier for the LLM and allows for a more straightforward comparison. The prompt fed to GPT-3.5-turbo was presented in Appendix Table 4. To evaluate the rationales, we randomly selected $50$ examples from each of the four datasets and asked GPT-3.5-turbo to determine rationales of which our method was better. We then aggregated the counts of "DSS is better," "Both are good," and "QCRD is better," as shown in Table 7. From the results, we observed that on SVAMP, CQA, and ESNLI, the model trained using our method generated better rationales than using DSS. However, on ANLI1, the model trained using DSS exhibited slightly better performance.
Table 7: The quality assessment results on the T5-base model for different sampling temperature settings, where three numbers represent counts of "DSS is better", "Both are good", and "QCRD is better", respectively.
@@ 5.4 Assessment for generated rationales
5.5 Distribution of rationale quality scores
The probability density estimation for the sampled rationale scores from the trained discriminator on the SVAMP test dataset is shown in Fig 6. Specifically, we considered the quality scores of: (1) positive and negative rationales from LLM’s sampled outputs (sampled 5 times); (2) negative rationales from sampling a trained T5-base model with temperature $\tau$ set to 1.5 and 2.0, respectively. It showed that the trained discriminator can effectively score different types of rationales. Scores of LLM’s positive rationales were around 0.95. For the trained student model, scores of the sampled negative rationales sometimes exceeded 0.7 (see the orange distribution), and it was necessary for the discriminator to assign low weights to these rationales. Furthermore, by comparing the orange distribution and the red distribution, we can see that the sampling temperature has a significant influence on the qualities of the rationales.
Figure 6: The probability density estimation for sampled rationale scores on the SVAMP test dataset, where rationales were from LLM’s and trained T5-base model’s sampled outputs, and $\tau$ denotes sampling temperature.
@@ 6 Conclusion
The knowledge distillation of CoT rationales from LLMs into smaller language models using a multi-task learning framework has been empirically shown to enhance performances of smaller language models. Building upon the framework, we introduces a general CoT distillation method, incorporating a contrastive learning perspective that considers both positive and negative knowledge. To generate positive and negative rationales, we propose an innovative approach that combines temperature sampling, the self-consistency of LLMs, and the self-adversarial of small language models themselves. Additionally, we develop a many-to-one contrastive distillation loss for better knowledge learning, where an online-update discriminator is used to judge qualities of rationales and assign them weights for optimizing the training process across the whole datasets. Extensive experiments conducted on multiple reasoning tasks demonstrate the superiority of our method over previous ones.
Limitations
Our work serves as a distillation method for language models, paving the way for further improvements. On one hand, as illustrated in Appendix A.5, it requires additional training time due to the distillation of sampled positive and online-inferenced negative rationales, even though our proposed method can enhance model performance without incurring additional deployment costs. On the other hand, the quality of knowledge for distillation is crucial. In this paper, we prompt the LLM to generate CoT rationales and further classify them into positive and negative ones through self-consistency. However, different types of prompts and decoding strategies can lead the LLM to produce various forms of positive CoT knowledge and more intuitive negative CoT knowledge, which may further improve the distillation effect.
@@ Appendix A Appendix
A.1 Details about Datasets
Following the setting in Hsieh et al. (2023), we provide detailed descriptions of the four benchmark datasets in Table 1. To illustrate the unbalanced proportion of positive and negative rationales from LLMs given the ground truth, we displayed the statistical description of the generated rationale annotations on training datasets for four benchmarks in Table 2. On the one hand, the number of positive rationales was larger than that of negative rationales (3.87:1.13). On the other hand, for many samples in the training dataset (more than $50\%$), there were only positive rationales. Therefore, there is a need to generate effective negative rationales in other ways.
@@ A.1 Details about Datasets
Table 2: Statistical descriptions of the generated rationale annotations, where r denotes rationale, and positive r achieves correct answers.
Dataset SVAMP CQA ANLI1 ESNLI
Average pos r (total 5) 3.87 3.89 3.93 3.31
Proportion with only pos r 0.55 0.68 0.66 0.50
Proportion with only neg r 0.08 0.13 0.11 0.20
A.2 Iteration-before-models for negative rationale generators
In this paper, we dynamically generated negative rationales using iteration-before-models through online temperature sampling. We took these iteration-before-models as negative generators, and we sampled them with a relatively high temperature value to generate negative rationales for every batch of datasets. As depicted in Fig. 1, to select a j-iteration-before-model for the negative rationale generator, we need to save a minimum of $j$ checkpoints for the model. This allows us to load the negative generator online and train the student model end-to-end instead of using a multi-turn approach. Additionally, as shown in Table 3, we found that the performance of the student models was not sensitive to the choice of $j$ from $\{3,5,10\}$, all of which outperformed the results obtained with a fixed negative generator (pretrained by DSS Hsieh et al. (2023)) or by using negative rationales derived from the self-consistency of the LLM.
Table 3: Results on SVAMP with different negative knowledge strategies, where the "Fixed" denotes fixing negative generator with the pretrained model, "SC" denotes using negative rationales from self-consistency.
Negative source $j$=3 $j$=5 $j$=10 Fixed SC
T5-base 68.0 69.0 68.5 66.5 64.5
Figure 1: A case of the j-iteration-before-model for the negative rationale generator.
Table 4: The prompt for GPT-3.5-turbo to judge rationales.
The prompt for GPT-3.5-turbo
A.3 Details about the Quality Judge
We incorporated a discriminator into our training process to assess the quality of rationales and assign corresponding weights to the losses. To construct the discriminator, we leveraged the encoder of the T5-base model along with one maxpooling layer and two linear layers to compute the quality score. Prior to training, the discriminator needs to be pretrained using the output rationales generated by LLMs with applying data augmentations to the negative rationales. Specifically, we employed word mask and replacement with the assistance of StanfordNLP Zeman et al. (2018) to balance the proportions of positive and negative rationales. The training objective is $L_{D}$ in (2). We pretrained the discriminator 500 max steps and we ensured scores for positive rationales close to $1$ and scores for negative rationales close to 00. The discriminator was further online-updated during training.
A.4 Influence of negative sampling temperature
The results of the T5-base model distilled by our method on SVAMP were displayed in Table 5 in terms of different negative sampling temperature settings. It was observed that when no sampling was performed (i.e., $\tau=0$) or a lower temperature value was used (i.e., $\tau=0.7$), the smaller model exhibited relatively poorer performance and showed larger fluctuations in accuracy. The best results were achieved when the temperature $\tau$ was set to $1.5$. The reason for this can be attributed to the fact that when the model approaches convergence, the output rationales with lower temperature values tend to be similar to the golden ones. Considering these similar outputs as negative samples can lead to detrimental effects.
Table 5: Results of our method on SVAMP with different sampling temperature $\tau$.
Temperature $\tau$ 0 0.7 1.5 2
T5-base 65.0 64.5 69.0 67.5
A.5 Computational cost
The training times for the T5-base and T5-small models using each method are shown in Table 6. Specifically, we trained the models on a single A100-80G GPU using the SVAMP benchmark. Compared to DSS Hsieh et al. (2023), our method requires an additional 9 hours for T5-base and 6 hours for T5-small. This increase is due to each input necessitating 5 positive rationales and 1 online-inferenced negative rationale for contrastive rationale distillation. However, we emphasize the motivation behind our method: to enhance the performance of deployed small language models. Our proposed QCRD effectively improves model performance without incurring additional parameter storage during deployment.
Table 6: The comparison for training time on T5-base and T5-small models, where D dnotes the Quality Judge.
Method Base/small training time (h)
Finetune 2.0 / 1.25
DSS 4.0 / 2.5
MI 4.2 / 2.6
QCRD 13.0 / 8.5
QCRD (w/o D) 12.0 / 7.5